At 25°C, K = 0.090 for the following reaction. H20(g) Cl20(g)2 HOCI(g) Calculate the concentrations of all species at equilibrium for each of the following cases. (a) 1.3 g H20 and 2.0 g Cl20 are mixed in a 1.1-L flask. [Cl20] [H2O] 四 (b) 1.5 mol pure HOCI is placed in a 1.7-L flask Cl20] [H2O]
Answer #1

(a)

1.3 g H2O (molar mass 18 g/mol) = \dfrac { \text { 1.3 g } }{ \text { 18 g/mol } } = \text { 0.0722 mol }

2.0 g Cl2O (molar mass 87 g/mol) \dfrac { \text { 2 g } }{ \text { 87 g/mol } } = \text { 0.02299 mol }

x moles of H2O will react with x moles of Cl2O to form 2x moles of HOCl.

The equilibrium concentrations are

[H_2O] = \frac{ \text { 0.0722 -x mol } }{ \text { 1.1 L } }

[Cl_2O] = \frac{ \text { 0.02299 -x mol } }{ \text { 1.1 L } }

[HOCl] = \frac{ \text { 2x mol } }{ \text { 1.1 L } }

K = \frac{[HOCl]^2}{[H_2O][Cl_2O]}

0.090 = \frac{[ \frac{ \text { 2x mol } }{ \text { 1.1 L } } ]^2}{ \frac{ \text { 0.0722 -x mol } }{ \text { 1.1 L } } \times \frac{ \text { 0.02299 -x mol } }{ \text { 1.1 L } } }

0.090 = \frac{ 4x^2}{ (0.0722 -x) \times ( 0.02299 -x )}

{ (0.0722 -x) \times ( 0.02299 -x )} = \frac{ 4x^2} { 0.090 }

{ (0.0722 -x) \times ( 0.02299 -x )} =44.44x^2

0.0722 ( 0.02299 -x ) -x( 0.02299 -x ) =44.44x^2

0.00166- 0.0722 x - 0.02299x+x^2=44.44x^2

43.44x^2+0.095211x-0.00166 =0

This is quadratic equation with solution

x = \frac{-b \pm \sqrt {b^2-4ac}}{2a}

x = \frac{-0.095211 \pm \sqrt {(0.095211)^2-4(43.44)(-0.00166)}}{2(43.44)}

x = \frac{-0.095211 \pm0.0545}{86.89}

x =0.00518 or x =-0.00737

The negative value is discarded since the number of moles cannot be negative.

x =0.00518

The equilibrium concentrations are

0.0722 -x mol 1.1 L 0.0722 -0.00518 mol 1H201 = = 0.061 M 1.1 L

[Cl_2O] = \frac{ \text { 0.02299 -x mol } }{ \text { 1.1 L } }= \frac{ \text { 0.02299 -0.00518 mol } }{ \text { 1.1 L } }=0.016 \: M

[HOCl] = \frac{ \text { 2x mol } }{ \text { 1.1 L } } = \frac{ \text { 2 (0.00518) mol } }{ \text { 1.1 L } }=0.0094 \: M

(b)

1.5 mol pure HOCl is placed in the flask

2x moles of HOCl will give x moles of H2O and x moles of Cl2O

The equilibrium concentrations are

[H_2O] = \frac{ \text { x mol } }{ \text { 1.7 L } }

[Cl_2O] = \frac{ \text { x mol } }{ \text { 1.7 L } }

[HOCl] = \frac{ \text { 1.5-2x mol } }{ \text { 1.7 L } }

K = \frac{[HOCl]^2}{[H_2O][Cl_2O]}

0.090 = \frac{[ \frac{ \text { 1.5-2x mol } }{ \text { 1.7 L } } ]^2}{ \frac{ \text {x mol } }{ \text { 1.7 L } } \times \frac{ \text { x mol } }{ \text { 1.7 L } } }

0.090 = \frac{ (1.5-2x)^2}{ (x) \times ( x )}

0.090 = \frac{ (1.5-2x)}{ x}

0.090 x= 1.5-2x

2.090 x= 1.5

x=0.718

The equilibrium concentrations are

[H_2O] = \frac{ \text { x mol } }{ \text { 1.7 L } }= \frac{ \text { 0.718 mol } }{ \text { 1.7 L } }=0.42 \: M

[Cl_2O] = \frac{ \text { x mol } }{ \text { 1.7 L } }= \frac{ \text { 0.718 mol } }{ \text { 1.7 L } }=0.42 \: M

[HOCl] = \frac{ \text { 1.5-2x mol } }{ \text { 1.7 L } }= \frac{ \text { 1.5-2(0.718) mol } }{ \text { 1.7 L } }=0.038 \: M